OpenStudy (calculusxy):

MEDAL FOR THE BEST ANSWER!!! Tom had 350 dollars to spend on 9 books. After buying them he had 17 dollars. How much did each book cost?

OpenStudy (anonymous):

350-17 = $333 Assuming he bought 9 books, $333/9 = $37/book However, it's unlikely that each book costed $37, some may cost more than others.

OpenStudy (anonymous):

We have no idea what each books cost. We only know that all 9 books cost $333.

OpenStudy (calculusxy):

My teacher told me that we were focusing on writing equations. We had to write an equation and then solve for the value of x.

OpenStudy (anonymous):

You cannot figure out the cost of each book. Unless you were told that each book had the same price. Otherwise, you can only write an equation for the total cost of the 9 books.

OpenStudy (calculusxy):

Yes, each book has the SAME cost.

OpenStudy (anonymous):

You didnt state that in the problem.

OpenStudy (calculusxy):

That wasn't written in the problem. But you can already tell that it has to be equal @Easyaspi314

OpenStudy (anonymous):

If it wasnt written, then you cannot assume they are equal. Sorry.

OpenStudy (calculusxy):

But please can you help me write an equation.

OpenStudy (calculusxy):

@Easyaspi314

OpenStudy (anonymous):

$350-9x = $17